# Spurious Correlations in Machine Learning: A Survey

### v1

@@ Abstract
Machine learning systems are known to be sensitive to spurious correlations between biased features of the inputs (e.g., background, texture, and secondary objects) and the corresponding labels. These features and their correlations with the labels are known as “spurious” because they tend to change with shifts in real-world data distributions, which can negatively impact the model’s generalization and robustness. In this survey, we provide a comprehensive review of this issue, along with a taxonomy of current state-of-the-art methods for addressing spurious correlations in machine learning models. Additionally, we summarize existing datasets, benchmarks, and metrics to aid future research. The paper concludes with a discussion of the recent advancements and future research challenges in this field, aiming to provide valuable insights for researchers in the related domains.
@@ 1 Introduction
While machine learning (ML) systems have achieved remarkable success, distribution shifts and adversarial examples have become critical issues Taori et al. (2020). It is becoming clearer that one of the major causes behind these issues is the high reliance on spurious correlations between biased features of the inputs (e.g., background, texture, and secondary objects) and the corresponding labels. When the correlations captured during training no longer hold in the test data, the performance of ML models tends to deteriorate, resulting in robustness issues and negative social impact in critical domains, such as healthcare. For example, previous trials in recognizing pneumonia with Convolutional Neural Networks (CNNs) show that these models rely on the metal tokens in the chest X-ray scans from different hospitals, instead of learning the pathological features of the disease Kirichenko et al. (2023). Therefore, it is critical to directly address the impact of spurious correlations.
In recent years, spurious correlations have been studied under various names, such as shortcuts, dataset biases, group robustness, simplicity bias, and so on. We have seen significant progress in analyzing and mitigating spurious correlations in various areas such as computer vision Wang et al. (2021), natural language processing Du et al. (2022b), and healthcare Huang et al. (2022). Despite the progress, there lacks a survey in this area that formally defines spurious correlations and comprehensively introduces and summarizes learning algorithms and other related problems to provide insights into future research.
@@ 2 Spurious Correlations
Spurious correlation, namely “correlations that do not imply causation” in statistics, refers to a situation where two variables appear to be related to each other, but the relationship is coincidental or confounded by an external variable. This phenomenon leads to misleading or incorrect interpretations of data and models. We first give a formal definition below.
Definition 1 (Spurious Correlation).
@@ 2.2 Why are Machine Learning Models Sensitive to Spurious Correlations?
2.2.1 Inductive Biases from Learning Algorithms
@@
2.2.2 Optimization
Formally, in the classical Empirical Risk Minimization (ERM) setting Vapnik (1991), let the triplet $(x,y,a)$ be a sample $x$ with its label $y$ and a spurious attribute $a$ which is typically unknown. Assume the group $g=(y,a)$ and $(x,y,a)$ follow the distributions $P_{g}$ and $P_{x|g}$ conditioned on $\mathcal{D}_{tr}$. Given a model with parameters $\theta$ as $f_{\theta}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ and a loss function $\ell:\mathbb{R}^{K}\times\mathcal{Y}\mapsto\mathbb{R}$, e.g., softmax cross-entropy, the objective is to minimize the training loss $\mathcal{L}_{\text{avg }}$:
@@ 2.2.2 Optimization
Since the optimization process does not take spurious attribute $a$ into consideration, when $y$ and $a$ exhibit a strong correlation, the ERM trained model will prioritize $a$ from $f_{\theta}(x)$ for predicting the label $y$ if $a$ is a relatively easier feature to learn. Upon testing on a dataset where this correlation is absent, the classifier’s reliance on the spurious attribute $a$ as a predictive feature can lead to incorrect predictions. It has been proved that machine learning models are optimized via ERM in Eq. 1 and have poor worst-group test errors Zhang et al. (2022) in Eq. 2.
@@
To achieve low worst-group test error, prior work Group DRO Sagawa et al. (2019) proposed modified objectives that focus on the worst-group loss which optimizes for the worst-group training loss or reweighting. However, there remains a strong assumption that group information ($\mathcal{G}$) is available during training. Without such information, different gradient descent algorithms may converge to different local minima. For example, gradient descent with large learning rates tends to learning simple patterns that are shared across examples, while it facilitates complex pattern learning and memorization with small learning rates Geirhos et al. (2020).
@@ 2.3 Related Areas
Domain Generalization. Domain generalization (DG), also known as out-of-distribution (OOD) generalization, aims to solve a problem where a model is trained on one or more related domains with the intent of generalizing to new, unseen test domains Wang et al. (2023). DG is a broad area of study and it varies depending on the assumptions made about the relationship between training and test distributions. A significant threat to domain generalization is spurious correlations, where features that are predictive in the training distribution no longer correlate with the label in the test distribution Yang et al. (2022).
Invariant Learning. The concept of invariant learning (IL) is creating models that utilize patterns or features in data that remain invariant across various domains Arjovsky et al. (2019); Chen et al. (2022). This is beneficial when training data originates from diverse sources, yet some fundamental principles or relationships are consistent. To accomplish invariant learning, it is crucial to train models in a way that they do not rely on spurious correlations.
Group Robustness. Group robustness refers to the consistency and fairness of a machine learning model’s performance across different groups or subpopulations within the data Liu et al. (2021); LaBonte et al. (2023). If a model depends on correlations unique to a specific group (for example, a correlation found in training data that predominantly represents a certain demographic), it may not perform well for other groups where these correlations don’t apply. One way to mitigate spurious correlation is from the group robustness perspective, aiming to maximize the accuracy of the worst-performing groups (usually the minority groups) while maintaining high average accuracy Yang et al. (2022).
Shortcut Learning. Shortcuts refer to spurious correlations that models learn, which perform well in standard benchmark environments but struggle in more complex testing domains Geirhos et al. (2020); Du et al. (2022a). This phenomenon, known as shortcut learning, often occurs when models are trained under empirical risk minimization (ERM). Such models find it difficult to generalize to relevant or challenging distribution shifts Yang et al. (2022).
Simplicity Bias. Simplicity bias (SB) is the inclination of deep neural networks to favor less complex patterns (such as the background) over more intricate ones (like the object’s semantics), even if all features have the same predictive power Tiwari and Shenoy (2023). This leads to vulnerabilities to minor adversarial perturbations (due to the small margin) and spurious correlations (with simpler features) Shah et al. (2020).
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7em,font=,, where level=2text width=11em,font=,, where level=3text width=em,font=,, where level=4text width=6.2em,font=,, [ Approaches to Address Spurious Correlation, ver [ Data Manipulation [ Data Augmentation, leaf ] [ Concept and Pseudo-label Discovery, leaf ] ] [ Representation Learning [ Causal Intervention, leaf ] [ Invariant Learning, leaf ] [ Feature Disentanglement, leaf ] [ Contrastive Learning, leaf ] ] [ Learning Strategy [ Optimization-based Methods, leaf ] [ Ensemble Learning, leaf ] [ Identification then Mitigation, leaf ] [ Finetuning Strategy, leaf ] [ Adversarial Training, leaf ] ] [ Other Methods ] ]
@@ 3.1.1 Data Augmentation
In some cases, data augmentation makes use of additional annotated data to leverage humans’ common sense knowledge of causality. One example of this is UV-DRO Srivastava et al. (2020), which uses human annotation to augment training examples with potential unmeasured variables. This reduces the spurious correlation problem to a covariate shift problem.
When additional annotation is not available, the following methods use existing training data to augment the dataset. Nuisances via Negativa Puli et al. (2022) proposes a data augmentation method to corrupt semantic information to produce models that can identify and adjust for where nuisances drive predictions. LISA Yao et al. (2022) is a simple mixup-based technique that learns invariant predictors via selective augmentation. This selectively interpolates samples either with the same labels but different domains or with the same domain but different labels. The Counterfactual Generator Zeng et al. (2020) creates counterfactual examples by making interventions on existing observational examples. This is used to enhance the original dataset and improve the performance of machine learning models.
3.1.2 Concept and Pseudo-label Discovery
Another line of work in data manipulation involves predefining concepts and generating pseudo-labels to enhance the conceptual information before training. This process helps break intrinsic spurious correlations within the training data and improve the model’s robustness. Discover and Cure (DISC) Wu et al. (2023) is a method that discovers unstable predefined concepts across different environments as spurious attributes. It then augments on the training data using the discovered concepts to reduce spurious correlations. Spread Spurious Attribute (SSA) Nam et al. (2022) is an example of generating pseudo-labels. It uses a small amount of spurious attribute annotated samples to estimate group information for the training set samples. The generated pseudo-labeled attribute together with the training set is then used to train a robust model using worst-case loss minimization.
@@ 3.2.1 Causal Intervention
Several methods consider the causal intervention as an implicit design inside the model structure. Wang et al. (2021) introduce a Causal Attention Module (CaaM) that utilizes unsupervised learning within the Vision Transformer architecture to self-annotate and mitigate confounding effects. By doing so, CaaM enhances model performance and promotes causality. Causal VQA Agarwal et al. (2020) uses a semantic editing-based approach to assess the robustness of Visual Question Answering (VQA) models. The model demonstrates significantly less flipping behavior under invariant and covariant semantic edits, resulting in improved causality. Counterfactual invariance measures how the predicted label behaves under interventions on portions of input data.
Other methods consider formal ways of adding structured causal graphs to reflect the true causal relationship between the attributes inside the data. Veitch et al. (2021) examines a true causal structure and applies causal inference tools to formalize and study perturbative stress tests. Latent Causal Invariance Models (LaCIM) Sun et al. (2021) represents a set of causal models that account for causal structure, aiding in the identification of causal relationships for predictions. De-confound-TDE Tang et al. (2020) offers a principled solution to pinpoint the causal effect of momentum in long-tailed classification tasks using causal graphs.
3.2.2 Invariant Learning
@@
Earlier work IRM Arjovsky et al. (2019) proposed an objective to learn deep, invariant features and correlations of data which are a complex function of latent variables, which has been a principled method in general domain generalization problems. It identifies and leverages features with consistent predictive performance, avoiding reliance on spurious correlations and ensuring stability across various domains.
@@
Several later works follow the path of IRM and try to emphasize different aspects of the invariant feature. From the risk perspective, REx Krueger et al. (2021) focuses on model performance across varied risk scenarios. By training models under different risk levels, REx encourages models to avoid overfitting to specific features of the training environment.
@@
From the environment perspective in the image data, EIIL Creager et al. (2021) aims to learn invariant features in the various environments of the images. It automates the inference of different environments within datasets and performs well especially when explicit labels of the environments are unavailable. SFB Eastwood et al. (2023) selectively amplifies features stable across environments, minimizing dependence on non-generalizable, spurious features, aiming to enhance the stability of features in the model.
@@
A more theoretical work Chen et al. (2022) analyzes the necessary conditions for group invariant learning for spurious-free predictors, and obtains two group criteria, i.e., falsity exposure and label balance. This paper thus proposes SCILL to satisfy the two criteria which is provably sufficient in solving spurious correlations.
3.2.3 Feature Disentanglement
@@
Feature Disentanglement method in machine learning aims to separate spurious and general representations within the latent space. These disentanglements are often achieved by a dual/multiple-branch design in the model structure. StableNet Zhang et al. (2021) gets rid of both linear and non-linear dependencies between features by utilizing the characteristics of Random Fourier Features (RFF), together with another standard classifier branch to compose up the final loss. Lee et al. (2021) trains two distinct encoders to separately extract bias features and intrinsic features. They further enhance model robustness by augmenting bias-conflicting samples through the random swapping of vectors between these encoders. Chroma-VAE Yang et al. (2022) utilizes a dual-pronged Variational Auto-Encoder (VAE) to disentangle latent representations and learn shortcut features in a smaller latent subspace. This method is focused on ensuring that these shortcut features do not dominate the learning process.
3.2.4 Contrastive Learning
@@
Contrastive learning has emerged as a promising approach in machine learning, particularly effective in addressing spurious correlations. This method was common in the self-supervised learning paradigm which can learn generalized representations. A notable implementation of this approach is Correct-n-Contrast (CnC) Zhang et al. (2022). CnC begins by training an ERM model to identify samples within the same class but with dissimilar spurious features. Following this, it employs a robust model trained through contrastive learning. This model is designed to learn similar representations for same-class samples, despite the presence of varying spurious features. The end result is a more generalizable and resilient model, capable of distinguishing between essential and non-essential features, thus mitigating the influence of spurious correlations.

### v2

@@ Abstract
Machine learning systems are known to be sensitive to spurious correlations between non-essential features of the inputs (e.g., background, texture, and secondary objects) and the corresponding labels. These features and their correlations with the labels are known as “spurious” because they tend to change with shifts in real-world data distributions, which can negatively impact the model’s generalization and robustness. In this paper, we provide a review of this issue, along with a taxonomy of current state-of-the-art methods for addressing spurious correlations in machine learning models. Additionally, we summarize existing datasets, benchmarks, and metrics to aid future research. The paper concludes with a discussion of the recent advancements and future challenges in this field, aiming to provide valuable insights for researchers in the related domains.
@@ 1 Introduction
While machine learning (ML) systems have achieved remarkable success, distribution shifts and adversarial examples have become critical issues (Taori et al., 2020). It is becoming clearer that one of the major causes behind these issues is the high reliance on spurious correlations between non-essential features of the inputs (e.g., background, texture, and secondary objects) and the corresponding labels. When the correlations captured during training no longer hold in the test data, the performance of ML models tends to deteriorate, resulting in robustness issues and negative social impact in critical domains, such as healthcare. For example, previous trials in recognizing pneumonia with Convolutional Neural Networks (CNNs) show that these models rely on the metal tokens in the chest X-ray scans from different hospitals, instead of learning the pathological features of the disease (Kirichenko et al., 2023). Therefore, it is critical to directly address the impact of spurious correlations.
In recent years, spurious correlations have been studied under various names, such as shortcuts, dataset biases, group robustness, simplicity bias, and so on. We have seen significant progress in analyzing and mitigating spurious correlations in various areas such as computer vision (Wang et al., 2021), natural language processing (Du et al., 2022b), and healthcare (Huang et al., 2022). Despite the progress, there lacks a survey in this area that formally defines spurious correlations and comprehensively introduces and summarizes learning algorithms and other related problems to provide insights into future research.
@@ 2 Spurious Correlations
Spurious correlation, namely “correlations that do not imply causation” in statistics, refers to a situation where two variables appear to be related to each other, but the relationship is coincidental or confounded by an external variable. This phenomenon leads to misleading or incorrect interpretations of data and models. We first give a formal definition.
Definition 2.1 (Spurious Correlation).
@@ 2.2 Why are Machine Learning Models Sensitive to Spurious Correlations?
Inductive Biases from Learning Algorithms
@@
Optimization
Formally, in the classical Empirical Risk Minimization (ERM) setting (Vapnik, 1991), let the triplet $(x,y,a)$ be a sample $x$ with its label $y$ and a spurious attribute $a$ which is typically unknown. Assume the group $g=(y,a)$ and $(x,y,a)$ follow the distributions $P_{g}$ and $P_{x|g}$ conditioned on $\mathcal{D}_{tr}$. Given a model with parameters $\theta$ as $f_{\theta}:\mathcal{X}\rightarrow\mathbb{R}^{K}$ and a loss function $\ell:\mathbb{R}^{K}\times\mathcal{Y}\mapsto\mathbb{R}$, e.g., softmax cross-entropy, the objective is to minimize the training loss $\mathcal{L}_{\text{avg }}$:
@@
Since the optimization process does not take spurious attribute $a$ into consideration, when $y$ and $a$ exhibit a strong correlation, the ERM trained model will prioritize $a$ from $f_{\theta}(x)$ for predicting the label $y$ if $a$ is a relatively easier feature to learn. Upon testing on a dataset where this correlation is absent, the classifier’s reliance on the spurious attribute $a$ as a predictive feature can lead to incorrect predictions. It has been proved that machine learning models are optimized via ERM in Eq. 1 and have poor worst-group test errors (Zhang et al., 2022) in Eq. 2.
@@
To achieve low worst-group test error, prior work Group DRO (Sagawa et al., 2019) proposed modified objectives that focus on the worst-group loss which optimizes for the worst-group training loss or reweighting. However, there remains a strong assumption that group information ($\mathcal{G}$) is available during training. Without such information, different gradient descent algorithms may converge to different local minima. For example, gradient descent with large learning rates tends to learning simple patterns that are shared across examples, while it facilitates complex pattern learning and memorization with small learning rates (Geirhos et al., 2020).
@@ 2.3 Related Areas
Domain Generalization. Domain generalization (DG), also known as out-of-distribution (OOD) generalization, aims to solve a problem where a model is trained on one or more related domains with the intent of generalizing to new, unseen test domains (Wang et al., 2023). DG is a broad area of study and it varies depending on the assumptions made about the relationship between training and test distributions. A significant threat to domain generalization is spurious correlations, where features that are predictive in the training distribution no longer correlate with the label in the test distribution (Yang et al., 2022).
Invariant Learning. The concept of invariant learning (IL) is creating models that utilize patterns or features in data that remain invariant across various domains (Arjovsky et al., 2019; Chen et al., 2022). This is beneficial when training data originates from diverse sources, yet some fundamental principles or relationships are consistent. To accomplish invariant learning, it is crucial to train models in a way that they do not rely on spurious correlations.
Group Robustness. Group robustness refers to the consistency and fairness of a machine learning model’s performance across different groups or subpopulations within the data (Liu et al., 2021; LaBonte et al., 2023). If a model depends on correlations unique to a specific group (for example, a correlation found in training data that predominantly represents a certain demographic), it may not perform well for other groups where these correlations don’t apply. One way to mitigate spurious correlation is from the group robustness perspective, aiming to maximize the accuracy of the worst-performing groups (usually the minority groups) while maintaining high average accuracy (Yang et al., 2022).
Shortcut Learning. Shortcuts refer to spurious correlations that models learn, which perform well in standard benchmark environments but struggle in more complex testing domains (Geirhos et al., 2020; Du et al., 2022a). This phenomenon, known as shortcut learning, often occurs when models are trained under empirical risk minimization (ERM). Such models find it difficult to generalize to relevant or challenging distribution shifts (Yang et al., 2022).
Simplicity Bias. Simplicity bias (SB) is the inclination of deep neural networks to favor less complex patterns (such as the background) over more intricate ones (like the object’s semantics), even if all features have the same predictive power (Tiwari & Shenoy, 2023). This leads to vulnerabilities to minor adversarial perturbations (due to the small margin) and spurious correlations (with simpler features) (Shah et al., 2020).
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=7em,font=,, where level=2text width=11em,font=,, where level=3text width=em,font=,, where level=4text width=6.2em,font=,, [ Methods to Address Spurious Correlation, ver [ Data Manipulation [ Data Augmentation, leaf ] [ Concept and Pseudo-label Discovery, leaf ] ] [ Representation Learning [ Causal Intervention, leaf ] [ Invariant Learning, leaf ] [ Feature Disentanglement, leaf ] [ Contrastive Learning, leaf ] ] [ Learning Strategy [ Optimization-based Methods, leaf ] [ Ensemble Learning, leaf ] [ Identification then Mitigation, leaf ] [ Finetuning Strategy, leaf ] [ Adversarial Training, leaf ] ] [ Other Methods ] ]
@@ 3.1 Data Manipulation
In some cases, data augmentation makes use of additional annotated data to leverage humans’ common sense knowledge of causality. One example of this is UV-DRO (Srivastava et al., 2020), which uses human annotation to augment training examples with potential unmeasured variables. This reduces the spurious correlation problem to a covariate shift problem.
When additional annotation is not available, the following methods use existing training data to augment the dataset. Nuisances via Negativa (Puli et al., 2022) proposes a data augmentation method to corrupt semantic information to produce models that can identify and adjust for where nuisances drive predictions. LISA (Yao et al., 2022) is a simple mixup-based technique that learns invariant predictors via selective augmentation. This selectively interpolates samples either with the same labels but different domains or with the same domain but different labels. The Counterfactual Generator (Zeng et al., 2020) creates counterfactual examples by making interventions on existing observational examples. This is used to enhance the original dataset and improve the performance of machine learning models.
Concept and Pseudo-label Discovery
Another line of work in data manipulation involves predefining concepts and generating pseudo-labels to enhance the conceptual information before training. This process helps break intrinsic spurious correlations within the training data and improve the model’s robustness. Discover and Cure (DISC) (Wu et al., 2023) is a method that discovers unstable predefined concepts across different environments as spurious attributes. It then augments on the training data using the discovered concepts to reduce spurious correlations. Spread Spurious Attribute (SSA) (Nam et al., 2022) is an example of generating pseudo-labels. It uses a small amount of spurious attribute annotated samples to estimate group information for the training set samples. The generated pseudo-labeled attribute together with the training set is then used to train a robust model using worst-case loss minimization.
@@ 3.2 Representation Learning
Several methods consider the causal intervention as an implicit design inside the model structure. (Wang et al., 2021) introduce a Causal Attention Module (CaaM) that utilizes unsupervised learning within the Vision Transformer architecture to self-annotate and mitigate confounding effects. By doing so, CaaM enhances model performance and promotes causality. Causal VQA (Agarwal et al., 2020) uses a semantic editing-based approach to assess the robustness of Visual Question Answering (VQA) models. The model demonstrates significantly less flipping behavior under invariant and covariant semantic edits, resulting in improved causality. Counterfactual invariance measures how the predicted label behaves under interventions on portions of input data.
Other methods consider formal ways of adding structured causal graphs to reflect the true causal relationship between the attributes inside the data. (Veitch et al., 2021) examines a true causal structure and applies causal inference tools to formalize and study perturbative stress tests. Latent Causal Invariance Models (LaCIM) (Sun et al., 2021) represents a set of causal models that account for causal structure, aiding in the identification of causal relationships for predictions. De-confound-TDE (Tang et al., 2020) offers a principled solution to pinpoint the causal effect of momentum in long-tailed classification tasks using causal graphs.
Invariant Learning
@@
Earlier work IRM (Arjovsky et al., 2019) proposed an objective to learn deep, invariant features and correlations of data which are a complex function of latent variables, which has been a principled method in general domain generalization problems. It identifies and leverages features with consistent predictive performance, avoiding reliance on spurious correlations and ensuring stability across various domains.
@@
Several later works follow the path of IRM and try to emphasize different aspects of the invariant feature. From the risk perspective, REx (Krueger et al., 2021) focuses on model performance across varied risk scenarios. By training models under different risk levels, REx encourages models to avoid overfitting to specific features of the training environment.
@@
From the environment perspective in the image data, EIIL (Creager et al., 2021) aims to learn invariant features in the various environments of the images. It automates the inference of different environments within datasets and performs well especially when explicit labels of the environments are unavailable. SFB (Eastwood et al., 2023) selectively amplifies features stable across environments, minimizing dependence on non-generalizable, spurious features, aiming to enhance the stability of features in the model.
@@
A more theoretical work (Chen et al., 2022) analyzes the necessary conditions for group invariant learning for spurious-free predictors, and obtains two group criteria, i.e., falsity exposure and label balance. This paper thus proposes SCILL to satisfy the two criteria which is provably sufficient in solving spurious correlations.
Feature Disentanglement
@@
Feature Disentanglement method in machine learning aims to separate spurious and general representations within the latent space. These disentanglements are often achieved by a dual/multiple-branch design in the model structure. StableNet (Zhang et al., 2021) gets rid of both linear and non-linear dependencies between features by utilizing the characteristics of Random Fourier Features (RFF), together with another standard classifier branch to compose up the final loss. (Lee et al., 2021) trains two distinct encoders to separately extract bias features and intrinsic features. They further enhance model robustness by augmenting bias-conflicting samples through the random swapping of vectors between these encoders. Chroma-VAE (Yang et al., 2022) utilizes a dual-pronged Variational Auto-Encoder (VAE) to disentangle latent representations and learn shortcut features in a smaller latent subspace. This method is focused on ensuring that these shortcut features do not dominate the learning process.
Contrastive Learning
@@
Contrastive learning has emerged as a promising approach in machine learning, particularly effective in addressing spurious correlations. This method was common in the self-supervised learning paradigm which can learn generalized representations. A notable implementation of this approach is Correct-n-Contrast (CnC) (Zhang et al., 2022). CnC begins by training an ERM model to identify samples within the same class but with dissimilar spurious features. Following this, it employs a robust model trained through contrastive learning. This model is designed to learn similar representations for same-class samples, despite the presence of varying spurious features. The end result is a more generalizable and resilient model, capable of distinguishing between essential and non-essential features, thus mitigating the influence of spurious correlations.